OpenStudy (anonymous):

given the quadratic function f(x) = x^2- x-6.determine the x-intercepts for y= f(2x)

OpenStudy (asnaseer):

firstly find the x-intercepts of f(x) - do you know how?

OpenStudy (anonymous):

yeah by factoring rite?

OpenStudy (asnaseer):

correct - so what do you get for this?

OpenStudy (anonymous):

(x-3) (x+2)

OpenStudy (asnaseer):

that is the correct factorisation - now do you know how to find the x-intercepts from this?

OpenStudy (anonymous):

umm..no

OpenStudy (anonymous):

oh wait, u mean 3 and -2?

OpenStudy (asnaseer):

think about what happens to the curve at the x-intercepts. what do you think will be the value of y when the curve touches the x-axis?

OpenStudy (asnaseer):

yes = you have the right values there :) now - do you know what to do next?

OpenStudy (anonymous):

nope :(

OpenStudy (asnaseer):

think about the graph of y=f(x) and that of y=f(2x)

OpenStudy (anonymous):

okay

OpenStudy (asnaseer):

for y=f(2x), the value of x would need to be halved in order to get the equivalent point on y=f(x)

OpenStudy (anonymous):

rite

OpenStudy (asnaseer):

therefore the x-intercepts for y=f(2x) will be half of the x-intercepts for y=f(x)

OpenStudy (asnaseer):

given that - what do you think will be the x-intercepts for y=f(2x)?

OpenStudy (anonymous):

isnt it zero? not sure

OpenStudy (asnaseer):

what values did you get for the x-intercepts for y=f(x) - look above

OpenStudy (anonymous):

oh 3 and -2

OpenStudy (anonymous):

so is it 1.5 and -1?

OpenStudy (asnaseer):

good - so what should the x-intercepts of y=f(2x) be?

OpenStudy (asnaseer):

thats it - well done! :)

OpenStudy (anonymous):

oh i get it know, so if it was say y=f(4x) rite what would have been the x interepts

OpenStudy (asnaseer):

what do you think?

OpenStudy (anonymous):

now*

OpenStudy (anonymous):

x intercepts would be 3/4 and -2/4

OpenStudy (asnaseer):

yes :)

OpenStudy (anonymous):

oh okay thanks soo much :)

OpenStudy (asnaseer):

yw :)